two parallel chords are drawn on the same side of the centre of circle of radius r. it is found that they subtend an angle of 2A and A at the centre of the circle. Then find the perpendicular distance between two chords

Grade:11

1 Answers

Sujit Kumar
6 years ago
y=r(Cos\frac{A}{2}-CosA)Sorry the last Step is wrong in the above solution. The answer is 
Let The perpendicular distance from center to side opposite to angle A be x
and the perpendicular distance from center to side opposite to angle 2A be x-y
Required to find:- The value of y
 
Applying Trigonometory,
 
On triangle with angle A
Cos\frac{A}{2}= \frac{x}{r}=>x=r(Cos\frac{A}{2})__________(1)
 
On triangle with angle 2A
CosA= \frac{x-y}{r}=>x-y=r(CosA)__________(2)
 
Subtracting Equation (2) from (1)
 
y=r(Cos\frac{A}{2}-CosA)
Ans: \ The \ perpendicular \ distance \ between \ the \ two \ parallel \ chords \ is \ r(Cos\frac{A}{2}-CosA)